OpenStudy (anonymous):

How do you do the math to get the answer to (x-h)(x^2+2xh+h^2)

OpenStudy (anonymous):

Are you solving for x?

OpenStudy (anonymous):

You want to factor the second term, (x^2 + 2xh + h^2) to make it (x + h)(x + h) Then assuming the equation is set equal to zero (x-h)(x^2+2xh+h^2) = 0 our three solutions for x are x - h = 0 x + h = 0 x + h = 0

OpenStudy (anonymous):

no i just want to foil the equation if you can please

OpenStudy (anonymous):

sorry i have been waiting so long I didnt realize you had replied

OpenStudy (anonymous):

:) no worries

OpenStudy (anonymous):

Remember that both of the terms in (x - h) are multiplied through. arranging it like this always helped me keep track of things better: (x-h)(x^2+2xh+h^2) = x (x^2+2xh+h^2) - h (x^2+2xh+h^2) Then it's easier to see exactly what you're multiplying. give it a shot :)

OpenStudy (anonymous):

cool are you sure 100%?

OpenStudy (anonymous):

That's not the complete foiled answer; I always thought it was easier to do the steps that way. But yes, I'm very, very sure this is correct up to here. Do you see why?